I have two questions.
1. Should the traingle be solved with Law of Sines or Laws of Cosines. Then solve the trangle and round to the nearest tenth. a=16, b=13, c=10
2. Should the triangle be solved beggining with Law of Sines or Laws of Cosines. Then solve the triangle and round to the nearest tenth. A=56 degrees, B=38 degrees, a-13.
Can some one please show me step by step how to do this. I have a test on it tonight and have to show my work.
Thanks

Answer by Alan3354(69443)

I have two questions.
1. Should the traingle be solved with Law of Sines or Laws of Cosines. Then solve the trangle and round to the nearest tenth. a=16, b=13, c=10
---------------
Since no angles are known, you have to use the law of cosines to find the first angle. Then you can use either to find the 2nd angle.
For angle A (opposite side a):
a^ = b^2 + c^2 - 2bc*cos(A)
256 = 169 + 100 - 260cos(A)
cos(A) = 13/260
A = 87.134º
The law so sines is quicker, so I'd use that for the other 2 angles.
sin(A)/a = sin(B)/b = sin(C)/c = 0.0624218
B = 54.24115º
C = 38.6248º
As a check, they add to 180º
--------------------------------
2. Should the triangle be solved beggining with Law of Sines or Laws of Cosines. Then solve the triangle and round to the nearest tenth. A=56 degrees, B=38 degrees, a-13.
--------------
2 angles are given, so the 3rd is 180-(56+38). C = 86.
Law of Sines is the only choice.
a/sin(56) = b/sin(38) = c/sin(86
13/sin(56) = 15.6808333
b = sin(38)*15.68... = 9.654
c = sin(86)*15.68... = 15.6426
-------------------------
Can some one please show me step by step how to do this. I have a test on it tonight and have to show my work.